# On the backward stability of s-step GMRES

### v1

@@ 3 An abstract framework for backward stability of $s$-step GMRES
Together with (5), (17), and Lemma 3, we have
$\begin{split}&\bigl{\|}M_{L}^{-1}(b-A\widehat{x}^{(i^{*})})\bigr{\rVert}\\ &\quad\leq\bigl{\|}M_{L}^{-1}b-M_{L}^{-1}A\widehat{Z}_{1:p}\widehat{y}^{(i^{*})}\bigr{\rVert}+\bigl{\|}M_{L}^{-1}A\Delta x^{(i^{*})}\bigr{\rVert}\\ &\quad\leq\bigg{\|}\begin{bmatrix}\widehat{r}&M_{L}^{-1}A\widehat{Z}_{1:p}\end{bmatrix}\begin{bmatrix}1\\ -\widehat{y}^{(i^{*})}\end{bmatrix}\bigg{\rVert}+\bigl{\|}\Delta r\bigr{\rVert}+\bigl{\|}M_{L}^{-1}\bigr{\rVert}\bigl{\|}A\bigr{\rVert}\bigl{\|}\Delta x^{(i^{*})}\bigr{\rVert}\\ &\quad\leq\bigg{\|}\bar{V}_{1:p+1}\widehat{R}_{1:p+1}(1)\begin{bmatrix}1\\ -\widehat{y}^{(i^{*})}\end{bmatrix}\bigg{\rVert}+\bigg{\|}\Delta F_{1:p+1}\begin{bmatrix}1\\ -\widehat{y}^{(i^{*})}\end{bmatrix}\bigg{\rVert}+\bigl{\|}\Delta r\bigr{\rVert}\\ &\qquad+\bigl{\|}M_{L}^{-1}\bigr{\rVert}\bigl{\|}A\bigr{\rVert}\bigl{\|}\Delta x^{(i^{*})}\bigr{\rVert}\\ &\quad\leq\sqrt{1+\omega_{p+1}}\,\bigg{\|}\widehat{R}_{1:p+1}(1)\begin{bmatrix}1\\ -\widehat{y}^{(i^{*})}\end{bmatrix}\bigg{\rVert}+\bigg{\|}\Delta F_{1:p+1}\begin{bmatrix}1\\ -\widehat{y}^{(i^{*})}\end{bmatrix}\bigg{\rVert}\\ &\qquad+\delta_{M_{L}^{-1}b}\bigl{\|}M_{L}^{-1}\bigr{\rVert}\bigl{\|}b\bigr{\rVert}+\varepsilon_{mv}\bigl{\|}M_{L}^{-1}\bigr{\rVert}\bigl{\|}A\bigr{\rVert}_{\mathsf{F}}\bigl{\|}\tilde{Z}_{1:p}\bigr{\rVert}_{\mathsf{F}}\bigl{\|}D_{1:p}\widehat{y}^{(i^{*})}\bigr{\rVert}\\ &\quad\leq\sqrt{1+\omega_{p+1}}\,(9\cdot\omega_{\kappa(r,W)}+10\cdot\omega_{\mathrm{qr}}+12\cdot\varepsilon_{ls})(1+2\cdot\delta_{M_{L}^{-1}b})\bigl{\|}M_{L}^{-1}\bigr{\rVert}\bigl{\|}b\bigr{\rVert}\\ &\qquad+\bigl{(}\sqrt{1+\omega_{p+1}}\,(1+\varepsilon_{AZ}+\delta_{M_{L}^{-1}AZ})\cdot(9\cdot\omega_{\kappa(r,W)}+10\cdot\omega_{\mathrm{qr}}+12\cdot\varepsilon_{ls}\\ &\qquad+12\cdot\omega_{H})+\varepsilon_{AZ}+\delta_{M_{L}^{-1}AZ}+\varepsilon_{mv}\bigr{)}\bigl{\|}M_{L}^{-1}\bigr{\rVert}\bigl{\|}A\bigr{\rVert}_{\mathsf{F}}\bigl{\|}\tilde{Z}_{1:p}\bigr{\rVert}_{\mathsf{F}}\bigl{\|}D_{1:p}\widehat{y}^{(i^{*})}\bigr{\rVert}.\end{split}$ (42)
@@
According to Remark 1, $\varepsilon_{r}$, $\varepsilon_{AZ}$, $\varepsilon_{ls}$, $\varepsilon_{mv}=O(\bm{u})$, and $\delta_{*}$ are highly preconditioner dependent. It remains to evaluate $1+\omega_{p+1}$, $\omega_{\mathrm{qr}}$, $\omega_{\kappa(r,W)}$, and $\omega_{H}$, which are all related only to the block orthogonalization method. Among these terms, it is usually easy to check that $1+\omega_{p+1}\leq O(\sqrt{n}\,)$ and $\omega_{\mathrm{qr}}=O(\bm{u})$ for different block orthogonalization methods.
@@ 7.1 Experiment settings
For constructing the linear systems $Ax=b$, we choose three sparse square matrices, shown in Table 1, from the SuiteSparse Matrix Collection [^1] https://sparse.tamu.edu and the Matrix Market [^2] https://math.nist.gov/MatrixMarket as $A$. The two with larger condition numbers from the Matrix Market are suggested as test problems in [6]. The other matrix from the SuiteSparse Matrix Collection is a relatively well-conditioned matrix. The right-hand vector $b$ is set to be the vector of all ones and the initial guess $x_{0}=0$.
@@ 7.2 Tests for different block size $s$
We illustrate how varying the block size $s$ impacts the relative backward error $\frac{\bigl{\|}Ax-b\bigr{\rVert}}{\bigl{\|}A\bigr{\rVert}_{\mathsf{F}}\bigl{\|}x\bigr{\rVert}+\bigl{\|}b\bigr{\rVert}}$ and the iteration count, as computed by the different $s$-step GMRES variants, in Figures 2, 3, and 4.
@@
Then we choose three specific values of $s$, i.e., $s=1$, $4$, $16$, to show the behavior of the backward error related to the iteration in Figures 5, 6, and 7. As Lemma 5 predicts, there is almost no chance to obtain a better solution after the “key dimension” is reached; Figures 57 illustrate that using the criterion (46) achieves the ”optimal” accuracy for most cases. Thus, it is necessary to employ (46) to test the “key dimension”.
@@ 8 Conclusions
We then provide an example, specifically Example 1, to illustrate the limitations of the classical $s$-step Arnoldi process (Algorithm 1). In certain scenarios, the condition number of the basis $B$ cannot be effectively managed by adaptively selecting $s$ to control the condition number of each sub-block of the basis, utilizing Newton or Chebyshev bases, or a restart process. Consequently, $s$-step GMRES may fail to achieve a satisfactory backward error even with a relatively small $s$. To address this issue, we introduce a modified $s$-step Arnoldi process that incorporates an extra QR factorization to make the basis $B$ well-conditioned. Numerical experiments demonstrate that this modified $s$-step Arnoldi process allows the use of a significantly larger $s$ while achieving the required accuracy. While the modified approach increases the communication cost per iteration by a factor of 2, it may be beneficial in cases where a higher $s$ can lead to greater per-iteration speedup. Future work involves high performance implementations on large-scale problems in order to better evaluate these tradeoffs.

### v2

@@ 3 An abstract framework for backward stability of $s$-step GMRES
Using (17) along with (7), $\bigl{\|}M_{L}^{-1}(b-A\widehat{x}^{(i^{*})})\bigr{\rVert}$ can be bounded by
$\begin{split}&\bigl{\|}M_{L}^{-1}(b-A\widehat{x}^{(i^{*})})\bigr{\rVert}\\ &\quad\leq\bigl{\|}M_{L}^{-1}b-M_{L}^{-1}A\widehat{Z}_{1:p}\widehat{y}^{(i^{*})}\bigr{\rVert}+\bigl{\|}M_{L}^{-1}A\Delta x^{(i^{*})}\bigr{\rVert}\\ &\quad\leq\bigg{\|}\begin{bmatrix}\widehat{r}&M_{L}^{-1}A\widehat{Z}_{1:p}\end{bmatrix}\begin{bmatrix}1\\ -\widehat{y}^{(i^{*})}\end{bmatrix}\bigg{\rVert}+\bigl{\|}\Delta r\bigr{\rVert}+\bigl{\|}M_{L}^{-1}\bigr{\rVert}\bigl{\|}A\bigr{\rVert}\bigl{\|}\Delta x^{(i^{*})}\bigr{\rVert}.\end{split}$
Furthermore, utilizing the bound of $\bigg{\|}\begin{bmatrix}\widehat{r}&M_{L}^{-1}A\widehat{Z}_{1:p}\end{bmatrix}\begin{bmatrix}1\\ -\widehat{y}^{(i^{*})}\end{bmatrix}\bigg{\rVert}$ derived from (52) and the bound of $\bar{V}_{1:p+1}$ from (12), we have
$\begin{split}&\bigl{\|}M_{L}^{-1}(b-A\widehat{x}^{(i^{*})})\bigr{\rVert}\\ &\quad\leq\bigg{\|}\bar{V}_{1:p+1}\widehat{R}_{1:p+1}(1)\begin{bmatrix}1\\ -\widehat{y}^{(i^{*})}\end{bmatrix}\bigg{\rVert}+\bigg{\|}\Delta F_{1:p+1}\begin{bmatrix}1\\ -\widehat{y}^{(i^{*})}\end{bmatrix}\bigg{\rVert}+\bigl{\|}\Delta r\bigr{\rVert}\\ &\qquad+\bigl{\|}M_{L}^{-1}\bigr{\rVert}\bigl{\|}A\bigr{\rVert}\bigl{\|}\Delta x^{(i^{*})}\bigr{\rVert}\\ &\quad\leq\sqrt{1+\omega_{p+1}}\,\bigg{\|}\widehat{R}_{1:p+1}(1)\begin{bmatrix}1\\ -\widehat{y}^{(i^{*})}\end{bmatrix}\bigg{\rVert}+\bigg{\|}\Delta F_{1:p+1}\begin{bmatrix}1\\ -\widehat{y}^{(i^{*})}\end{bmatrix}\bigg{\rVert}\\ &\qquad+\delta_{M_{L}^{-1}b}\bigl{\|}M_{L}^{-1}\bigr{\rVert}\bigl{\|}b\bigr{\rVert}+\varepsilon_{mv}\bigl{\|}M_{L}^{-1}\bigr{\rVert}\bigl{\|}A\bigr{\rVert}_{\mathsf{F}}\bigl{\|}\tilde{Z}_{1:p}\bigr{\rVert}_{\mathsf{F}}\bigl{\|}D_{1:p}\widehat{y}^{(i^{*})}\bigr{\rVert},\end{split}$
which further implies, by (53) and Lemma 3,
$\begin{split}&\bigl{\|}M_{L}^{-1}(b-A\widehat{x}^{(i^{*})})\bigr{\rVert}\\ &\quad\leq\sqrt{1+\omega_{p+1}}\,(9\cdot\omega_{\kappa(r,W)}+10\cdot\omega_{\mathrm{qr}}+12\cdot\varepsilon_{ls})(1+2\cdot\delta_{M_{L}^{-1}b})\bigl{\|}M_{L}^{-1}\bigr{\rVert}\bigl{\|}b\bigr{\rVert}\\ &\qquad+\bigl{(}\sqrt{1+\omega_{p+1}}\,(1+\varepsilon_{AZ}+\delta_{M_{L}^{-1}AZ})\cdot(9\cdot\omega_{\kappa(r,W)}+10\cdot\omega_{\mathrm{qr}}+12\cdot\varepsilon_{ls}\\ &\qquad+12\cdot\omega_{H})+\varepsilon_{AZ}+\delta_{M_{L}^{-1}AZ}+\varepsilon_{mv}\bigr{)}\bigl{\|}M_{L}^{-1}\bigr{\rVert}\bigl{\|}A\bigr{\rVert}_{\mathsf{F}}\bigl{\|}\tilde{Z}_{1:p}\bigr{\rVert}_{\mathsf{F}}\bigl{\|}D_{1:p}\widehat{y}^{(i^{*})}\bigr{\rVert}.\end{split}$ (54)
@@
According to Remark 1, $\varepsilon_{r}$, $\varepsilon_{AZ}$, $\varepsilon_{ls}$, $\varepsilon_{mv}=O(\bm{u})$, and the $\delta_{*}$ terms are highly preconditioner dependent. It remains to evaluate $1+\omega_{p+1}$, $\omega_{\mathrm{qr}}$, $\omega_{\kappa(W)}$, $\omega_{\kappa(r,W)}$, and $\omega_{H}$, which are all related only to the block orthogonalization method. Among these terms, it is usually easy to check that $1+\omega_{p+1}\leq O(\sqrt{n}\,)$ and $\omega_{\mathrm{qr}}=O(\bm{u})$ for different block orthogonalization methods. In addition, $\omega_{\kappa(W)}$ can be obtained from the connection with $\omega_{\kappa(r,W)}$ as shown in Lemma 3. Determining $\omega_{\kappa(r,W)}$ is crucial for the backward error, since Theorem 1 demonstrates that $\omega_{\kappa(r,W)}$ directly affects the backward error, while $\omega_{\kappa(W)}$ determines the possibility of achieving the backward error specified in equation (3.38) for the given linear system.
@@ 7.1 Experiment settings
For constructing the linear systems $Ax=b$, we choose three commonly-used sparse square matrices, shown in Table 2, from the SuiteSparse Matrix Collection [^1] https://sparse.tamu.edu and the Matrix Market [^2] https://math.nist.gov/MatrixMarket as $A$. The two with larger condition numbers from the Matrix Market are suggested as test problems in [6]. The other matrix from the SuiteSparse Matrix Collection is a relatively well-conditioned matrix. The right-hand vector $b$ is set to be the vector of all ones and the initial guess $x_{0}=0$. We have selected these particular linear systems for demonstration purposes because the standard GMRES algorithm can converge relatively quickly even without the use of a preconditioner.
@@ 7.2 Tests for different block size $s$
We illustrate how varying the block size $s$ impacts the relative backward error $\frac{\bigl{\|}Ax-b\bigr{\rVert}}{\bigl{\|}A\bigr{\rVert}_{\mathsf{F}}\bigl{\|}x\bigr{\rVert}+\bigl{\|}b\bigr{\rVert}}$ and the iteration count (in which each $s$-step counts as s iterations) for the different $s$-step GMRES variants in Figures 2, 3, and 4.
@@
In Figures 2, 3, and 4, notice that the number of iterations suddenly decreases when using the additional criteria (58). To help clarify this observation, we then choose three specific values of $s$, i.e., $s=1$, $4$, $16$, to show the behavior of the backward error related to the iteration in Figures 5, 6, and 7. As Lemma 5 predicts, there is almost no chance to obtain a better solution after the “key dimension” is reached; Figures 57 illustrate that using the criterion (58) achieves the ”optimal” accuracy for most cases. Thus, it is necessary to employ (58) to test the “key dimension”.
@@ 8 Conclusions
We then provide an example, specifically Example 1, to illustrate the limitations of the classical $s$-step Arnoldi process (Algorithm 1). In certain scenarios, the condition number of the basis $B$ cannot be effectively managed by adaptively selecting $s$ to control the condition number of each sub-block of the basis, utilizing Newton or Chebyshev bases, or a restart process. Consequently, $s$-step GMRES may fail to achieve a satisfactory backward error even with a relatively small $s$. To address this issue, we introduce a modified $s$-step Arnoldi process that incorporates an extra QR factorization to make the basis $B$ well-conditioned. Numerical experiments demonstrate that this modified $s$-step Arnoldi process allows for the use of a significantly larger $s$ while achieving the required accuracy. While the modified approach increases the communication cost per iteration by a factor of 2, it may be beneficial in cases where a higher $s$ can lead to greater per-iteration speedup. Future work involves high performance implementations on large-scale problems in order to better evaluate these tradeoffs.